OpenStudy (anonymous):

Please Help.. c^2=10^2+6^2

jimthompson5910 (jim_thompson5910):

\[\large c^2=10^2+6^2\] \[\large c^2=100+36\] \[\large c^2=136\] \[\large c=??\]

OpenStudy (anonymous):

What are the ?? marks stand for?

jimthompson5910 (jim_thompson5910):

they are for you to fill in the blank

jimthompson5910 (jim_thompson5910):

for instance, if we had 2x = 10 and I said x = ?? then you would fill in x = 5

OpenStudy (anonymous):

so c=68?

jimthompson5910 (jim_thompson5910):

c^2 = 136 doesn't mean 2 times c = 136

jimthompson5910 (jim_thompson5910):

c^2 means c squared

jimthompson5910 (jim_thompson5910):

so how do you undo the square to isolate c?

OpenStudy (anonymous):

You divide?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

thats if we had 2c = 136

jimthompson5910 (jim_thompson5910):

to undo the square, we take the square root of both sides

jimthompson5910 (jim_thompson5910):

So \[\large c^2=136\] \[\large c=\sqrt{136}\] Now we simplify the root to get \[\large c=\sqrt{4*34}\] \[\large c=\sqrt{4}*\sqrt{34}\] \[\large c=2\sqrt{34}\]

OpenStudy (anonymous):

and thats it?

jimthompson5910 (jim_thompson5910):

yep